could someone check to see if I have done this right. subtract 7/6 - (-13/18
7/6 - (-13/18)= 21/18- (-13/18) = 21/18 - 13/18= 8/18=-8/18

Answer by jim_thompson5910(35256)

When you subtract a negative, you're adding a positive.


So 21/18- (-13/18) = 21/18 + 13/18 = (21+13)/18 = 34/18


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim